If a balloon has a volume of 1.329 L at a pressure of 2.265 atmospheres, what will be the volume of the balloon if the pressure is increased to 4.993 atmospheres?

Give your answer to four significant figures.
Answer #1
  1. for a given quantity of gas pressure in inversly proportional to volume px 1 (Boyles law) so Ple=k (k = constant) this Piul
Answer #2

SOLUTION :


As per Boyle’s Law,


P V = constant


So,


P1 V1 = P2 V2

=> V2 = (P1/P2) * V1 

=> V2 = (2.265 / 4.993) * 1.329 

=> V2 = 0.6029 L 


So, volume of balloon at increased pressure of 4.993 atm  will be 0.6029 L (ANSWER).


answered by: Tulsiram Garg